OpenStudy (anonymous):

help and explain your answer For the investment with interest compounded annually, find the final balance using the formula A = P(1 + r)t $36,000 for 4.5 years at 3% A. $37,080.00 B. $40,860.00 C. $41,121.60 D. $117,232.38

OpenStudy (anonymous):

@Puahina please help me

OpenStudy (anonymous):

Just a minute....

OpenStudy (anonymous):

ok and thanx for the medal

OpenStudy (paki):

by using the formula... A = P(1 + r)*t here p= $36,000 time= 4.5 years rate= 3% putting all values in equation... we get... A=36000*(1+3%)^4.5 A=36000*(1+3/100)^4.5... take lcm inside the brackets here... ok... A=36000*(100+3/100)^4.5 A=36000*(103/100)^4.5 A=36000*(1.03)^4.5.... now resolve the issue of power here... A=36000*(1.142266687) A=41121.60$... @lululife

OpenStudy (anonymous):

@paki u answered my question so late i already took the test and i failed sadly and im retaking it so thanx anyways

OpenStudy (paki):

oh.... it mean you were cheating... :-) by the way, i was no calculator and found it very late... and best of luck... @lululife

OpenStudy (anonymous):

what im not cheating i needing help on how to do this and you can take the test three times its diffrent quetions '.....' how am i cheating i wanted someone to help me on that question but no one helped me in like26 minutes so i got 2 wrong and failed and im re taking it why are u smiling when u said i cheated its a litttle creepy ......... @Paki

OpenStudy (paki):

@lululife oh dear, i was just kidding... grow up... if you mind it, then sorry...

OpenStudy (anonymous):

didnt seem like it and sorry i took it so seriously im kinda childish sorry again @paki

OpenStudy (paki):

ahhh its ok.... keep smiling.... @lululife